# D-LiFT: Improving LLM-based Decompiler Backend via Code Quality-driven Fine-tuning

### v1

@@ 4.2.3 Readability Metric
To evaluate the readability of each candidate refinement, $o_{i}$, fairly, we generate a relative readability score by directly comparing it against the original decompiled code, $o_{g}$. To do this, we customize established software-engineering readability measures by incorporating selected R2I features as follows:
@@
To compute $R_{b\&w}(o_{i},o_{g})$, we follows the B&W framework, which defines a feature set, $\textbf{f}=\{f_{1},f_{2},...\}$, e.g., average number of commas per line, and their associated weights, $\textbf{w}=\{w_{1},w_{2},...\}$. Specifically, we first apply B&W to $o_{i}$ and $o_{g}$ to obtain two absolute scores. We then compute their relative difference and pass this value through a sigmoid function to map it into the range [-1,1], yielding a normalized, relative readability score. Formally,
@@
$\text{where}\quad R_{mul}(x)=(\sum\textbf{f}(x)\cdot\textbf{w})$
Nevertheless, certain general features, e.g., total line number, remain outside the scope of $R_{b\&w}(o_{i},o_{g})$. Fortunately, $R_{R2I}(o_{i},o_{g})$ includes these features, f and their associated weights w within its conflicting features and generic features categories. Our implementation follows the methodology established in the original research paper to generate the $R_{R2I}(o_{i},o_{g})$. Through renormalization of the feature-associated weights, $R_{R2I}(o_{i},o_{g})$ produces values within the range [-1,1], calculated using the following approach:
$R_{R2I}$ $(o_{i},o_{g})=\sum\textbf{w}\cdot r_{elog}(\textbf{f}(o_{i})-\textbf{f}(o_{g})),$ (11)
@@
For $\gamma$ and $\delta$ in Equation 9, we allow users to adjust their values according to their specific needs.
5 Implementation
We implemented D-LiFT as a modular Python framework, using the state-of-the-art open-source decompiler, Ghidra (version 11.2).
Reinforcement learning. We train our models using the TRL library [59] (version 0.18) from Huggingface. During the RL, we set num_generations to 3, num_iterations to 10, per_device_train_batch_size to 3, vllm_gpu_memory_utilization to 0.7, leaving all other parameters at their default values.
Semantic check-related settings. For the Recompiler, we follow its default configuration by setting the maximum iteration count to 10.
For the semantics checking, we implement our semantic check framework on angr [61] (version 9.2), z3 [14] (version 4.13), and prompt [69] (verison 1.0). To reduce the training time, we set the execution timeout to 30 seconds.
D-Score Metric settings. We set $syn_{pen}$ to -3, $ret_{pen}$ to -2, and $call_{pen}$ to -1.5. For readability metric, we set $\gamma$ to 0.25 and $\delta$ to 0.75 for Equation 9, which makes $R_{read}(o_{i},o_{g})\in(-1,1)$. These values follow the Equation 4 and experimental observations to prevent sparse rewards [49].
6 Evaluation
@@
  • In Section 6.1, we describe our experimental setup, including the evaluation of the reward metric and function (i.e., D-Score), dataset details, and baseline model selection.
  • In Section 6.2 and Section 6.3, we demonstrate how effectively D-LiFT enhances LLM-generated decompiled code via the above setup.
  • In Section 6.4, we provide concrete case studies illustrating how D-LiFT refines specific decompiled snippets to improve both accuracy and readability.
6.1 Experiment Setup
In this section, we begin by evaluating D-Score along two dimensions, applicability and precision, in Section 6.1.1. We then describe our training/evaluation dataset in Section 6.1.2. Finally, we detail our baseline model selections and reasons behind them in Section 6.1.3.
6.1.1 D-Score Evaluation
We conduct experiments on a cluster node equipped with an NVIDIA A100 Tensor Core GPU (80GB of memory) [43], two 32-core AMD EPYC 7543 CPUs, and 400 GB of RAM.
Dataset. Our metric evaluation uses a dataset of 1,948 decompiled functions sourced from binaries that previous literature uses, including coreutils [13] (v9.5) and util-linux [58] (v2.41). Specifically, we first use GCC [1] to compile these projects with -O2 optimization under x86 Linux platform and then ran Ghidra [3] (version 11.2) on 578 resulting binaries and object files, yielding 5,385 unique decompiled functions (1,792 from coreutils and 3,593 from util-linux). To focus on functions with sufficient complexity and room for improvement, we then filtered this set to include only those with at least 20 lines of code and a cyclomatic complexity [41] greater than 3, leaving 1,948 (36.2%) functions (653 from coreutils and 1,295 from util-linux).
Methodology. D-Score contains two core components, accuracy check and readability assessment. Because its readability metric is an aggregation of two proven metrics [10, 17], our evaluation focuses on the decompiled code accuracy examination. To do this, we prompt Qwen-Coder-2.5-3B [6] with the above 1,948 decompiled functions and then pass each of its generated code completions, along with the corresponding original binary, through D-Score for scoring. We evaluate D-Score along two dimensions: (1) Applicability: The proportion of functions for which D-Score can successfully derive the symbolic models from the binary. (2) Precision: Among these analyzable functions, the fraction for which D-Score correctly determines that the generated code is semantically equivalent to the original.
@@ 3 D-Score errors 14%
TABLE II: The accuracy, precision, recall, and F1 score of D-Score on the tested decompiler.
@@
Applicability result. Our evaluation shows that out of 1,968 functions in our evaluation program set, D-Score can test 79.9% (1,573) of these functions, i.e., generate the symbolic models of these functions from the binary without errors. To understand the limitation, we randomly select 50 functions where D-Score fails and manually analyze them. Table II categorizes the failures encountered by D-Score and reports their occurrence frequency. These errors fall into three categories: (1) Errors from the underlying tool, e.g., bugs in angr or incorrect identification of function boundaries when encountering no-return calls, (2) Timeout due to the scale of a function, and (3) Internal errors of D-Score, e.g., unsupported floating-point instructions.
@@
Precision result. We evaluate the precision of D-Score by randomly sampling 100 functions from the 1,573 that D-Score can analyze and manually compare each LLM-generated output against its original binary’s semantics to verify whether the decision made by D-Score was correct. Table II presents our results in terms of accuracy, precision, recall, and F1 score. Our analysis uncovers that: 1 False positives occur because the memory model of D-Score allocates fresh memory addresses for symbolized pointers. Even when the generated code is semantically correct, the values of pointers differ from those in the original binary, causing the symbolic executor to report a spurious mismatch. (2) False negatives occur because we assume all external function calls return zero. When using a default return value in a conditional instruction, one branch will never be explored during the symbolic execution of the binary. If the LLM’s output omits code for those unvisited branches, D-Score cannot detect the discrepancy and thus overlooks the error.
6.1.2 Dataset
We begin by briefly analyzing the accuracy of the data that D-Score can process, and then explain how we construct our training and evaluation datasets.
@@ 6.1.2 Dataset
Dataset accuracy analysis. As detailed in Section 6.1.1, since D-Score is only able to work on 1,573 functions, we base our training and evaluation datasets on this subset. To better understand our dataset, we first analyze the score distribution from D-Score over these 1,573 decompiled outputs generated directly from the decompiler. Running D-Score on the Ghidra-generated functions allows us to partition them into two groups: originally accurate (i.e., error-free) and originally inaccurate. Of the 1,573 functions, 1,025 are classified as originally accurate (i.e., D-Score = 0, since $R_{read}(o_{g},o_{g})$ = 0), while the remaining 548 fall into originally inaccurate (i.e., D-Score < 0), 307 exhibiting syntax errors and 241 containing semantic errors.
@@
Training and evaluation dataset. We randomly choose 300 functions from the originally accurate group as the training data. This selection criterion is based on the fact that LLMs demonstrate a limited ability to correct decompiler-introduced inaccuracies when provided only with decompiled code, as listed in Section 6.3. To train the model, we use the following template: “prompt”: [ { “role”: “system”, “content”: “You are a helpful assistant for improving the decompiled result from the user. The user will input the decompiled result from Ghidra. Please improve its readability while preserving its semantics. Please do not add comments. Please just output the improved code.” }, { “role”: “user”, “content”: “[original decompiled code]” } .
@@
We evaluate D-LiFT on the remaining 1,273 functions, which we split into two subsets: 725 functions that are originally accurate (OA) and 548 functions that are originally inaccurate (OIA).
@@
TABLE III: Different baseline models’ performance on the originally accurate (OA) decompiled code. Specifically, it shows how many functions become inaccurate after being processed by the baseline model (i.e., LLM without fine-tuning).
6.1.3 Baseline Model
Due to computational resource constraints, we restrict our LLM to fewer than 3 billion parameters. Specifically, as an RL framework, GRPO is particularly memory-intensive, since it requires simultaneous inference to generate candidate outputs and backpropagation to update model parameters, both of which consume significant GPU memory. Additionally, reward normalization in GRPO mandates generating at least two candidates per input, further increasing memory demands. In our experiments, even after using vLLM [30] with a memory-efficient setting of 0.7, we still encountered out-of-memory errors when running D-LiFT on models larger than 3 billion parameters.
Therefore, we select two code-focused LLMs, Qwen2.5-Coder [26] (1.5B and 3B variants) and one general-purpose model, Llama3.2-3B [21]. We excluded the prior work LLM4Decompile-End-1.3B [55] from our main evaluation after observing its poor performance on our dataset. Specifically, we ask all four above models to improve our 1,025 originally accurate functions, score the outputs using D-Score and show the results in Table III. Since being fine-tuned based on the model released from November 2023, LLM4Decompile-End-1.3B improved only 70 functions (6.8%), whereas 955 functions (93.2%) turned inaccurate. In contrast, the three newer LLMs (released around September 2024) made only 44.4% of functions inaccurate. Given this marked difference, we omit LLM4Decompile-End-1.3B from further analysis.
@@ 6.1.3 Baseline Model
6.2 D-LiFT Performance On OA
We show the evaluation of D-LiFT on originally accurate functions (OA) in two steps. In Section 6.2.1, we compare outputs from the baseline and fine-tuned models using D-Score. Then Section 6.2.2 analyzes the underlying reasons for the differences. Finally, in Section 6.2.3, we show how LLMs’ performance is affected by the size of the training dataset.
6.2.1 Result
Figure 5 summarizes the number of original decompiled functions within the OA dataset that are improved by the baseline and fine-tuned models and their average D-Score. To this end, we compare the output from each model with the original decompiled code. For each model, D-LiFT improves the quality of 55.3% more functions on average, compared to the baseline LLMs. By selecting the best output among all three baseline models, 509 (70.2%) functions (avg. +0.465) show improvement, while fine-tuned models achieve improvement in 616 functions (84.9%) averaging +0.517.
Figure 6: Overall performance of D-LiFT on the OA dataset. This chart shows, for each function, which source, baseline LLM, fine-tuned LLM, or original decompiler achieves the highest D-Score score and reports the percentage of functions in each category.
To better quantify the improvements brought by fine-tuning, we compared the output of each fine-tuned model against its corresponding baseline and the original decompiled output. To do this, for each function, we select the version with the highest D-Score score among the three and show the result in Figure 6. Specifically, for Qwen2.5-Coder-1.5B, fine-tuned output is best for 355 (49.0 %) functions (avg. +0.399), baseline LLM for 159 (21.9 %) functions (avg. +0.419). For Qwen2.5-Coder-3B, fine-tuned output is best for 361 (49.8%) functions (avg. +0.402) and baseline LLM for 130 (17.9%) functions (avg. +0.398). For Llama3.2-3B, fine-tuned output is best for 312 (43.0%) functions (avg. +0.523) and baseline LLM for 165 (22.8%) functions (avg. +0.561). Moreover, by taking the best output among all six LLMs and the original decompiled code, 625 (86.2 %) functions showed improvement (average +0.585), while only 100 functions remained unimproved. The above result shows that D-LiFT helps the fine-tuned LLMs dominate their performance in generating high-quality decompiled code.
@@ 6.2.1 Result
TABLE IV: This table shows how many functions get improved and regressed by D-LiFT. This comparison happens between the output of the fine-tuned model and the baseline model, where the improvement refers to the function get higher score from the fine-tuned model and the regression refers to the function get higher score from the baseline model.
@@
6.2.2 Findings

### v2

@@ IV-B3 Readability Metric
To evaluate the readability of the candidate code, $o_{i}$, fairly, we generate a relative readability score by directly comparing it against the original decompiled code, $o_{g}$. However, as described in Section˜III-B, none of the existing ones can be directly applied in our context. To address this limitation, we customize established software engineering readability models by integrating selected features from the only decompiler-related metric, R2I [17], and the most classic readability metric, B&W [10] as follows:
@@
, where $\gamma$ and $\delta$ weight each metric’s contribution. Our metric adopts the same principle as R2I and B&W, aggregating multiple code features by multiplying each feature’s count by its weight and summing the results. Moreover, to blend them effectively, we normalize their individual scores to a fixed interval, allowing precise control over their relative influence.
To compute $R_{b\&w}(o_{i},o_{g})$, we follow the B&W framework, which defines a feature set, $\textbf{f}_{b\&w}=\{f_{1},f_{2},$...\}, e.g., average number of commas per line, and their associated weights, $\textbf{w}_{b\&w}=\{w_{1},w_{2},$...\}. Specifically, we first apply B&W to $o_{i}$ and $o_{g}$ to obtain two absolute scores. We then compute their relative difference and pass this value through a sigmoid function to map it into the range [-1,1], yielding a normalized, relative readability score. Formally,
@@
$\text{where}\quad R_{mul}(x)=(\sum\textbf{f}_{b\&w}(x)\cdot\textbf{w}_{b\&w})$
Nevertheless, certain decompilation-specific features remain outside the scope of $R_{b\&w}(o_{i},o_{g})$. Fortunately, $R_{R2I}(o_{i},o_{g})$ includes these features, $\textbf{f}_{R2I}$ and their associated weights $\textbf{w}_{R2I}$. Our implementation follows the methodology established in the original research paper to generate the $R_{R2I}(o_{i},o_{g})$. Through renormalization of the feature-associated weights, $R_{R2I}(o_{i},o_{g})$ produces values within the range [-1,1], calculated using the following approach:
$R_{R2I}$ $(o_{i},o_{g})=\sum\textbf{w}_{R2I}\cdot r_{elog}(\textbf{f}_{R2I}(o_{i})-\textbf{f}_{R2I}(o_{g})),$ (11)
@@
V Implementation
We implemented D-LiFT as a modular Python framework, using the SOTA open-source decompiler, Ghidra (version 11.2).
V-A Reinforcement learning
We run TRL [60] (version 0.18) from Huggingface to train models. We set num_generations to 3, num_iterations to 10, per_device_train_batch_size to 3, vllm_gpu_memory_utilization to 0.7, leaving all other parameters as default.
V-B Semantic check-related settings
For the Recompiler, we follow its default configuration by setting the maximum iteration count to 10.
For the semantics checking, we implement our semantic check framework on angr [62] (version 9.2), z3 [14] (version 4.13), and prompt [71] (version 1.0). To reduce the training time, we set the execution timeout to 30 seconds.
V-C D-Score Metric settings
We set $syn_{pen}$ to -3, $ret_{pen}$ to -2, and $call_{pen}$ to -1.5. For readability metric, we set $\gamma$ to 0.25 and $\delta$ to 0.75 for Equation˜9, making $R_{read}(o_{i},o_{g})\in(-1,1)$. These values follow Equation˜4 and experimental observations to prevent sparse rewards [50]. We leverage the fact that both B&W and R21 are defined and validated via user studies [17, 10], reflecting key readability features of decompiled code such as the number of references, do-while constructs, and goto labels.
VI Evaluation
@@
  • In Section˜VI-A, we describe our experimental setup, including the evaluation of D-Score, dataset details, and baseline model selection.
  • In Section˜VI-B and Section˜VI-C, we demonstrate how effectively D-LiFT enhances LLM-generated decompiled code via the above setup.
  • In Section˜VI-D, we provide concrete case studies illustrating how D-LiFT refines specific decompiled snippets to improve both accuracy and readability.
VI-A Experiment Setup
In this section, we begin by evaluating D-Score along two dimensions, applicability and precision, in Section˜VI-A1. After that, we describe our training and evaluation dataset in Section˜VI-A2. Finally, we detail our baseline model selections and reasons behind them in Section˜VI-A3.
VI-A1 D-Score Evaluation
We conduct experiments on a cluster node with an NVIDIA A100 Tensor Core GPU (80GB) [43], two 32-core AMD EPYC 7543 CPUs, and 400 GB of RAM.
Dataset. Our metric evaluation uses a dataset of 1,948 decompiled functions sourced from binaries that the most classic previous literature uses, which are coreutils [13] (v9.5) and util-linux [59] (v2.41). Specifically, we first use GCC [1] to compile these projects with -O2 optimization under x86 Linux platform and then run Ghidra [3] (version 11.2) on 578 resulting binaries and object files, yielding 5,385 unique decompiled functions (1,792 from coreutils and 3,593 from util-linux). To focus on functions with sufficient complexity and room for improvement, we then filter this set to include only those with at least 20 lines of code and a cyclomatic complexity [41] greater than 3, resulting in 1,948 (36.2%) functions (653 from coreutils and 1,295 from util-linux).
Methodology. D-Score contains two core components, accuracy check and readability assessment. Because our readability metric is an aggregation of two established metrics [10, 17], we accept its validity. Hence, our validity evaluation focuses on the metrics of decompiled code accuracy. To do this, we prompt Qwen-Coder-2.5-3B [6] with the above 1,948 decompiled functions and then pass each of its generated code completions, along with the corresponding original binary, through D-Score for an accuracy check. Specifically, we evaluate the accuracy check of D-Score along two dimensions: (1) Applicability: The proportion of functions for which D-Score can successfully derive the symbolic models from the binary. (2) Precision: Among these analyzable functions, the fraction for which D-Score correctly determines that the generated code is semantically equivalent to the original.
@@ 3 D-Score errors 14%
TABLE II: The accuracy, precision, recall, and F1 of D-Score on LLM-generated output.
@@
Applicability result. D-Score successfully finishes the accuracy check of 80.7% (1,573/1,948) of functions, i.e., generates the symbolic models of these functions and gets the result from the SMT solver without errors. To understand the limitation, we randomly sample 50 functions, where D-Score fails, without replacement, and manually analyze them. Table˜II categorizes the failures encountered by D-Score and reports their occurrence frequency. These errors fall into three categories: (1) Errors from the underlying tool, e.g., bugs in angr or incorrect identification of function boundaries when encountering no-return calls, (2) Timeout due to the scale of a function, and (3) Internal errors of D-Score, e.g., unsupported floating-point instructions.
@@
Precision result. We evaluate the precision of D-Score by randomly sampling 100 functions without replacement from the 1,573 that D-Score can analyze, and manually compare each LLM-generated output against its original binary’s semantics to verify whether the decision made by D-Score was correct. Table˜II presents our results in terms of accuracy, precision, recall, and F1 score. Our analysis uncovers that: (1) False positives occur because the memory model of D-Score allocates fresh memory addresses for symbolized pointers. Even when the generated code is semantically correct, the values of pointers differ from those in the original binary, causing the symbolic executor to report a spurious mismatch. (2) False negatives occur because we assume all external function calls return a constant value. When using a constant return value in a conditional instruction, one branch may never be explored during the symbolic execution of the binary. However, since users can vary the assumed constant, rerunning D-Score with multiple return values may exercise most branches and thus eliminate these false negatives. If the LLM’s output omits code for those unvisited branches, D-Score cannot detect the discrepancy and thus overlooks the error.
VI-A2 Dataset
We first present an overview of the dataset, then explain how we construct our training and evaluation datasets.
@@ VI-A2 Dataset
Dataset overview. Given that D-Score supports only 1,573 functions (see Section˜VI-A1), we center our overview on this filtered set. To better understand our dataset, we first run D-Score on the raw Ghidra outputs of these functions and examine their results distribution. Based on the accuracy check, the functions were split into two groups: originally accurate (i.e., error-free) and originally inaccurate. Of the 1,573 functions, 1,025 are classified as originally accurate (i.e., pass D-Score’s accuracy check), while the remaining 548 fall into originally inaccurate (i.e., fail the D-Score’s accuracy check), 307 exhibiting syntax errors and 241 containing semantic errors.
@@
Training and evaluation. We randomly choose 300 functions from the originally accurate group as the training data. This selection criterion is based on the fact that LLMs demonstrate a limited ability to correct decompiler-introduced inaccuracies when provided only with decompiled code, as listed in Section˜VI-C. We apply this template uniformly during both training and inference: “prompt”: [ { “role”: “system”, “content”: “You are a helpful assistant for improving the decompiled result from the user. The user will input the decompiled result from Ghidra. Please improve its readability while preserving its semantics. Please do not add comments. Please just output the improved code.” }, {“role”: “user”, “content”: “[original decompiled code]”}.
@@
We evaluate D-LiFT on the remaining 1,273 functions with two subsets: 725 functions that are originally accurate (OA) and 548 functions that are originally inaccurate (OIA).
@@
TABLE III: Different baseline models’ performance on the originally accurate (OA) decompiled code. Specifically, it shows how many functions become inaccurate after being processed by the baseline LLM (i.e., without fine-tuning).
VI-A3 Baseline Model
Due to computational resource constraints, we restrict our LLM to fewer than 3 billion parameters. Specifically, as a reinforcement learning framework, GRPO is particularly memory-intensive, since it requires simultaneous inference to generate candidate outputs and backpropagation to update model parameters, both of which consume significant GPU memory. Additionally, the reward normalization in GRPO mandates generating at least two candidates per input, further increasing memory demands. In our experiments, even after using vLLM [30] with a memory-efficient setting of 0.7, we still encountered out-of-memory errors when running D-LiFT on models larger than 3 billion parameters.
Therefore, we select two code-focused LLMs, Qwen2.5-Coder [26] (1.5B and 3B variants) and one general-purpose model, Llama3.2-3B [21]. As shown in Table˜III, due to the poor performance of LLM4Decompile-End-1.3B [56] on our dataset, where 93% of the functions go wrong after applying it, we excluded it from our main evaluation. More specifically, we ask all four above models to improve our 1,025 originally accurate functions, score the outputs using D-Score (i.e., -3 if the syntax check is not passed, -2 or -1.5 if the semantic check is not passed and score in (-1,1) for the readability) and show the results in Table˜III. Since being fine-tuned based on the model released from November 2023, LLM4Decompile-End-1.3B made 955 functions (93.2%) inaccurate. In contrast, the three newer LLMs (released around September 2024) made only 44.4% of functions inaccurate. Given this marked difference, we omit LLM4Decompile-End-1.3B from further analysis.
@@ VI-A3 Baseline Model
VI-B D-LiFT Performance on OA
We show the evaluation of D-LiFT on originally accurate functions (OA) in two steps. We demonstrate changes brought from D-LiFT in Section˜VI-B1. Section˜VI-B2 analyzes the underlying reasons for the differences.
VI-B1 Result
To clearly present, we split the result into two parts: (1) a comparison between baseline and fine-tuned LLMs, and (2) an assessment of D-LiFT ’s overall performance.
Fine-tuned LLMs vs. baseline LLMs. Figure˜6 summarizes the number of original decompiled functions within the OA dataset that are improved by the baseline and fine-tuned models and their average D-Score. To this end, we compare the output from each model with the original decompiled code. Compared with the baseline LLMs without D-Score-driven fine-tuning, our trained LLMs on average improve 55.3% more decompiled functions. By selecting the best output among all three baseline LLMs, 509 (70.2%) functions (avg. +0.465) show improvement, while fine-tuned LLMs achieve improvement in 616 functions (84.9%), averaging +0.517.
Figure 7: Overall performance of the selection system within D-LiFT on the OA dataset. This chart shows, for each function, which source, baseline LLM, fine-tuned LLM, or original decompiler achieves the highest D-Score score and reports the percentage of functions in each category.
@@ VI-B1 Result
TABLE IV: This table shows how many functions get improved and regressed by the training of D-LiFT. This comparison happens between the output of the fine-tuned model and the baseline model, where the improvement refers to the function getting a higher score from the fine-tuned model and the regression refers to the function getting a higher score from the baseline model.
@@
Overall improvement by D-LiFT. Figure˜7 shows the result of the selection system (i.e., selecting the version with the highest D-Score score among the fine-tuned model, the baseline model, and the original decompiled output). Specifically, for Qwen2.5-Coder-1.5B, fine-tuned LLM’s output is the best for 355 (49.0 %) functions (avg. +0.399), baseline LLM for 159 (21.9 %) functions (avg. +0.419). For Qwen2.5-Coder-3B, fine-tuned LLM’s output is best for 361 (49.8%) functions (avg. +0.402) and baseline LLM for 130 (17.9%) functions (avg. +0.398). For Llama3.2-3B, fine-tuned LLM’s output is best for 312 (43.0%) functions (avg. +0.523) and baseline LLM for 165 (22.8%) functions (avg. +0.561). Moreover, by taking the best output among all six LLMs and the original decompiled code, 625 (86.2%) functions showed improvement (average +0.585), while only 100 functions remained unimproved.
VI-B2 Findings